Solve using Matrices-----
4x-y=5
3x-3y=0

Found 2 solutions by TimothyLamb, Alan3354:
Answer by TimothyLamb(4379)

---
4x-y=5
3x-3y=0
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
system matrix: A
4 -1
3 -3
---
unknown vector: X
x
y
---
constant vector: B
5
0
---
inverse of system matrix: Ai
0.33333333 -0.11111111
0.33333333 -0.44444444
---
determinant of system matrix = -9
---
matrix equation:
X = Ai * B
---
solution (X vector):
x = 1.6666667
y = 1.6666667
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php

Answer by Alan3354(69443)

Solve using Matrices-----
4x-y=5
3x-3y=0
------------
|4 -1|
|3 -3|
Det = 4*-3 - 3*-1 = -12 + 3 = -9
---
For x:
|-1 -5|
|-3 0|
x*Det = -1*0 - (-3*-5) = -15
x = -15/-9 = 5/3
------
For -y*Det:
|4 -5|
|3 0|
-y*Det = 4*0 - 3*-5 = 15
-y = 15/-9
y = 5/3